i have an example problem that looks like this..
(X-12)^2=16
^2..means to the second power.
please help me understand this.

Answer by stanbon(57387)

(X-12)^2=16
---
(x-12)^2-16 = 0
Factor:
[(x-12)-4][(x-12)+4] = 0
---
Simplify:
[x-16][x-8] = 0
---
Solve:
x = 16 or x = 8
=======================
Cheers,
Stan H.